# Functional Latent Dynamics for Irregularly Sampled Time Series Forecasting

### v1

@@ 1 Introduction
Figure 1: Example for regularly and irregularly sampled Time Series with three channels. The observations and forecasting targets are marked as black crosses.
@@
  • We provide a Proof-of-Concept on a simple toy datasets task that is generated with the Goodwin oscillator model [4], an ODE designed to model enzyme synthesis.
@@
Our code is publicly available on an anonymous Git repository: https://anonymous.4open.science/r/Functional-Latent_Dynamics-5B1E/
@@ 2 Problem Formulation
An Irregularly sampled multivariate time series (IMTS) is a sequence $x:=\left(\left(t_{1},v_{1}\right),\ldots,\left(t_{N},v_{N}\right)\right)$ of $N$ many pairs where each pair consists of an observation time point $t_{n}\in\mathbb{R}$ and observation event $v_{n}\in X^{C}:=(\mathbb{R}\,\cup\,\{\text{NaN}\})^{C}$ made at $t_{n}$; $C\in\mathbb{N}$ is the number of channels, $v_{n,c}\neq\text{NaN}$ represents a observed value and $v_{n,c}=\text{NaN}$ represents a missing value. An IMTS forecasting query is a sequence $t^{q}:=(t^{q}_{1},\ldots,t^{q}_{K})$ of time points for which observation values are sought (where $\min_{k=1:K}\,t^{q}_{k}>\max_{n=1:N}\,t_{n}$). Any sequence $y:=(y_{1},\ldots,y_{K})$ of same length with values in $X^{C}$ we call an IMTS forecasting answer. To measure the difference between the ground truth forecasting answer $y$ (possibly with missing values) and the predicted forecasting answer $\hat{y}$ (without missing values), a scalar loss function $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ such as squared error is averaged over all query time points and non-missing observations:
@@
where $N_{k}=|\{c\in[C]\mid y_{k,c}\neq\text{NaN}\}|$ denotes the amount of non-missing values of the forecasting answer $y_{k}$ at time point $t^{k}_{q}$.
@@ 3 Background
Here, $\tau$ can both reference to observation time points $t$ and query time points $t^{q}$. $f$ is a trainable neural network that governs the dynamics of the hidden state. The hidden state $z(\tau)$ is defined at all times and can be evaluated at any desired time-points. Hence, they are a natural fit to model IMTS, where observation times are continuous. However, a numerical ODE solver is required to infer the hidden state:
@@
GRU-ODE-Bayes [3] integrates a continuous version of Gated Recurrent Units (GRU) into the neural ODE architecture and updates $z(t)$ with Bayesian inference based.
@@
While ODE-based models have the advantage of learning from continuous time observations, they require complex numerical ODE solver which is slow [1]. Additionally, they process the observations of an IMTS sequentially worsening the run time and also increase the memory requirements. Furthermore, ODE-based models cannot directly handle missing values. Typically, they require missing value indicators which act as additional input channels in the series complicating the learning process.
@@ 4 Functional Latent Dynamics
Inferring the hidden state at any time point with a function is computationally efficient if that function is simple and does not depend on other time points. Large portion of the literature applies sequential models such as RNNs to learn the inductive bias from the causal nature of time series. However, they are slow, as they have to operate sequentially. Alternatively, recent transformer based works in similar domains [10, 15] show that we can achieve state-of-the-art performance even without applying the sequential model. Hence, in this work, we use simple curves such as polynomial (linear in eq. 3, quadratic in eq. 4) or sine functions (in eq. 5).
@@
Here, sin is applied coordinate-wise. Once we compute the latent state $z$, we apply a multilayer feedforward neural network ($\text{NN}^{\text{out}}$) to compute $\hat{y}$ via
@@ 5 Inferring Coefficients
$\phi^{h}_{d}(t):=\begin{cases}a_{di}t+b_{di}\quad&\text{if }d=1\\ \sin(a_{di}t+b_{di})\quad&\text{if }1<d\leq D\end{cases}$ (6)
Here, $a_{di}$ and $b_{di}$ are trainable parameters. This embedding helps to learn periodic terms from the sinusoidal embeddings and non-periodic terms from the linear embedding [13].
@@ 7.2 Competing models
We compare FLD models against members of the neural ODE family: GRU-ODE-Bayes [3], Neural Flows [1], LinODENet [12], CRU [11]. Besides the ODE-based models, we also compare our results to GraFITi [17], the state-of-the-art in IMTS forecasting model.
@@ 7.5 Results
Table 3, Table 4 and Table 5 show the Test MSEs for each model on the 75%-3, 75%-25% and 50%-50% task respectively. Based on our results we do not observe an FLD variant which consistently outperforms the other two members of the model family. For most datasets, FLD-L shows to be the best fit for the short and medium forecasting range, while FLD-S has the best accuracy on two datasets for the 50%-50% task among FLD variants. FLD-Q makes the best predictions on USHCN for the 75%-3 task, where it even surpasses the state-of-the art model GraFITi [17]. However, USHCN carries large standard deviations across all models and task, especially for the longer forecasting ranges. Consequently, findings on this dataset are less conclusive. GraFITi reports superior forecasting accuracy’s on 10 out twelve dataset/task combinations, but on Physionet-2012 and the 75%-25% task FLD-L improves on GraFITi making it the state-of-the art in this part of the evaluation. When we compare FLD’s performance to the ODE-based models, we observe that the most accurate FLD variant outperforms the best ODE-based models in 7 out of 12 cases. In particular, LinODENet outperforms, FLD on all datasets for the 50%-50% task.

### v2

@@ 1 Introduction
Figure 1: Example for regularly and irregularly sampled Time Series with two channels. The observations and forecasting targets are marked as black crosses.
@@
  • We provide a Proof-of-Concept on a simple toy dataset that is generated with the Goodwin oscillator model [4], an ODE designed to model enzyme synthesis.
@@
Our code is publicly available on an anonymous Git repository: https://github.com/kloetergensc/Functional-Latent_Dynamics
@@ 2 Problem Formulation
An Irregularly sampled multivariate time series (IMTS) is a sequence $x:=\left(\left(t_{1},v_{1}\right),\ldots,\left(t_{N},v_{N}\right)\right)$ of $N$ many pairs where each pair consists of an observation time point $t_{n}\in\mathbb{R}$ and observation event $v_{n}\in X^{C}:=(\mathbb{R}\,\cup\,\{\text{NaN}\})^{C}$ made at $t_{n}$; $C\in\mathbb{N}$ is the number of channels, $v_{n,c}\neq\text{NaN}$ represents an observed value and $v_{n,c}=\text{NaN}$ represents a missing value. An IMTS forecasting query is a sequence $t^{q}:=(t^{q}_{1},\ldots,t^{q}_{K})$ of time points for which observation values are sought (where $\min_{k=1:K}\,t^{q}_{k}>\max_{n=1:N}\,t_{n}$). Any sequence $y:=(y_{1},\ldots,y_{K})$ of same length with values in $X^{C}$ we call an IMTS forecasting answer. To measure the difference between the ground truth forecasting answer $y$ (possibly with missing values) and the predicted forecasting answer $\hat{y}$ (without missing values), a scalar loss function $\ell:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ such as squared error is averaged over all query time points and non-missing observations:
@@
where $N_{k}=|\{c\in[C]\mid y_{k,c}\neq\text{NaN}\}|$ denotes the amount of non-missing values of the forecasting answer $y_{k}$ at time point $t^{q}_{k}$.
@@ 3 Background
Here, $\tau$ can both reference to observation time points $t$ and query time points $t^{q}$. $f$ is a trainable neural network that governs the dynamics of the hidden state. The hidden state $z(\tau)$ is defined and can be evaluated at any desired time-point. Hence, they are a natural fit to model IMTS, where observation times are continuous. However, a numerical ODE solver is required to infer the hidden state:
@@
GRU-ODE-Bayes [3] integrates a continuous version of Gated Recurrent Units (GRU) into the neural ODE architecture and updates $z(t)$ with Bayesian inference.
@@
While ODE-based models have the advantage of learning from continuous time observations, they require a complex numerical ODE solver which is slow [1]. Additionally, they process the observations of an IMTS sequentially worsening the run time and also increase the memory requirements. Furthermore, ODE-based models cannot directly handle missing values. Typically, they require missing value indicators which act as additional input channels in the series complicating the learning process.
@@ 4 Functional Latent Dynamics
Inferring the hidden state at any time point with a function is computationally efficient if that function is simple and does not depend on other time points. Large portion of the literature applies sequential models such as RNNs to learn the inductive bias from the causal nature of time series. However, they are slow, as they have to operate sequentially. Alternatively, recent transformer based works in similar domains [10, 15] show that we can achieve state-of-the-art performance even without applying the sequential model. Hence, in this work, we use simple curves such as polynomial (linear (FLD-L) in eq. 3, quadratic (FLD-Q) in eq. 4) or sine (FLD-S) functions (in eq. 5).
@@
Here, sin is applied coordinate-wise. Once we have computed the latent state $z$, we apply a multilayer feedforward neural network ($\text{NN}^{\text{out}}$) to compute $\hat{y}$ via
@@ 5 Inferring Coefficients
$\phi^{h}_{d}(t):=\begin{cases}a_{dh}t+b_{dh}\quad&\text{if }d=1\\ \sin(a_{dh}t+b_{dh})\quad&\text{if }1<d\leq D\end{cases}$ (6)
Here, $a_{dh}$ and $b_{dh}$ are trainable parameters. This embedding helps to learn periodic terms from the sinusoidal embeddings and non-periodic terms from the linear embedding [13].
@@ 7.2 Competing Models
We compare FLD models against members of the neural ODE family: GRU-ODE-Bayes [3], Neural Flows [1], LinODENet [12], CRU [11]. Besides the ODE-based models, we also compare our results to GraFITi [17], the state-of-the-art in IMTS forecasting.
@@ 7.5 Results
Table 3, Table 4 and Table 5 show the Test MSEs for each model on the 75%-3, 75%-25% and 50%-50% task respectively. Based on our results we do not observe an FLD variant which consistently outperforms the other two members of the model family. For most datasets, FLD-L shows to be the best fit for the short and medium forecasting range, while FLD-S has the best accuracy on two datasets for the 50%-50% task among FLD variants. FLD-Q makes the best predictions on USHCN for the 75%-3 task, where it even surpasses the state-of-the art model GraFITi [17]. However, USHCN carries large standard deviations across all models and task, especially for the longer forecasting ranges. Consequently, findings on this dataset are less conclusive. GraFITi reports superior forecasting accuracy’s on 10 out 12 dataset/task combinations, but on Physionet-2012 and the 75%-25% task FLD-L improves on GraFITi making it the state-of-the art in this part of the evaluation. When we compare FLD’s performance to the ODE-based models, we observe that the most accurate FLD variant outperforms the best ODE-based models in 7 out of 12 cases. In particular, LinODENet outperforms, FLD on all datasets for the 50%-50% task.